# MIA-Bench: Towards Better Instruction Following Evaluation of Multimodal LLMs

### v3

@@ Abstract
We introduce MIA-Bench, a new benchmark designed to evaluate multimodal large language models (MLLMs) on their ability to strictly adhere to complex instructions. Our benchmark comprises a diverse set of 400 image-prompt pairs, each crafted to challenge the models’ compliance with layered instructions in generating accurate responses that satisfy specific requested patterns. Evaluation results from a wide array of state-of-the-art MLLMs reveal significant variations in performance, highlighting areas for improvement in instruction fidelity. Additionally, we create extra training data and explore supervised fine-tuning to enhance the models’ ability to strictly follow instructions without compromising performance on other tasks. We hope this benchmark not only serves as a tool for measuring MLLM adherence to instructions, but also guides future developments in MLLM training methods.
Figure 1: An example from MIA-Bench, featuring an image and a complex instruction to test models’ compliance with layered instructions that are compositional in nature. Responses from GPT-4v [1] and InternVL-v1.5 [2] are evaluated using GPT-4o as the judge.
@@ 1 Introduction
Figure 2: Comparison of various multimodal LLM benchmarks. (Left) Fixed-form visual question answering, often features short answers or multi-choice formats, such as MMMU [3]. This format is popular due to its ease of evaluation. (Middle) Open-ended responses, such as LLaVA-Bench (in the Wild) [4]. (Right) The proposed MIA-Bench, which also uses open-ended responses but focuses on evaluating precise adherence to complex instructions within the prompt.
The rapid advancement of Multimodal Large Language Models (MLLMs) [1, 4, 5, 6, 7, 8, 9, 10, 11] has been a defining feature of recent AI research, showcasing increased model capabilities to comprehend and respond to visual inputs, often termed as multimodal “instruction following”.
To measure the progress of instruction following, many multimodal benchmarks have been developed, which can be roughly divided into two broad categories: ($i$) fixed-form visual question answering (VQA), often with short answers or using a multi-choice QA format; and ($ii$) free-form conversations with open-ended responses. Many current benchmarks have adopted the first format, including VQAv2 [12], TextVQA [13], ScienceQA [14], MME [15], MMBench [16], SEED-Bench [17], MathVista [18], and MMMU [3]. These benchmarks are popular due to their ease of use in evaluating metrics and presenting model comparisons.
However, as visual assistant models, the ability to engage users in free-form conversations is also crucial. Benchmarks in this format include LLaVA-Bench [4], MM-Vet [19], VisIT-Bench [20], InfiMM-Eval [21], and the most recent Vibe-Eval [22] and LLaVA-Bench-Wilder [23]. Typically, the free-form model responses are evaluated using external models as the judge. These benchmarks are closer to daily-life visual chat scenarios; however, the type of “instruction following” examined in these benchmarks usually gauges a model’s ability to perform tasks in a broad, often loosely defined manner. Yet, the precise adherence to complex instructions within prompts – a critical aspect for evaluating LLMs [24, 25, 26] – remains less explored in the context of multimodal LLMs.
@@
We evaluate a wide array of MLLMs on the proposed benchmark, ranging from closed-source models (e.g., GPT-4o [27], Gemini Pro [10], Claude-3 [28], Reka [29]) to open-source ones (e.g., LLaVA-NeXT [30], Intern-VL-Chat-1.5 [2], CogVLM2 [8], Phi-3-Vision [31]). Our investigations reveal notable variations in model performance, highlighting great opportunities for improvement.
@@
Our contributions are summarized as follows. ($i$) We construct MIA-Bench, a new benchmark to comprehensively evaluate MLLMs on their capability to strictly adhere to instructions. ($ii$) We provide a detailed analysis of popular MLLMs, and suggest training methods for enhanced instruction following. For this purpose, we created training data and conducted experiments for additional supervised fine-tuning. MIA-Bench will be open-sourced, and we hope this benchmark can serve as a useful resource to stimulate further research on multimodal instruction adherence.
@@ 2 MIA-Bench
MIA-Bench consists of 400 image-prompt pairs, with examples shown in Figure 3. The images are collected from diverse sources, including COCO 2017 validation set [32], SBU [33], TextVQA [34], and Flickr. Images in the Flickr subset are photos of a variety of themes, including animals, art, architectures, text, food, math, etc. Images from the other three sources are randomly sampled from each corresponding source. Figure 4 shows the top 15 image content categories and the distribution of the 8 sub-instruction categories in MIA-Bench. The image content is labeled using GPT-4v. For each image, we manually write diverse and challenging instructions that contain multiple sub-instructions.
@@
  • Image-dependent. MMStar [35] pointed out that on some multimodal benchmarks, MLLMs can generate correct answers without accessing images half of the time. Multi-modal capabilities are necessary to correctly answer MIA-Bench prompts.
@@ 2.1 Instruction Categories
Figure 5 shows the most frequently used verbs and co-occurring nouns in MIA-Bench. To guarantee the diversity of prompts, when writing the instructions, we contribute instructions of various levels of complexity: basic, intermediate, advanced, creative, and complex. The basic category is the simplest; the instructions normally only contain one or two sub-instructions, such as “What is the color of the cat?”, or “Describe the sofa in two words.”. The intermediate category consists of instructions that contain three or more sub-instructions, but are in general easy for MLLMs to follow. The advanced category contains instructions that are challenging and contain three or more sub-instructions. The creative category contains instructions that instruct MLLMs to generate creative pieces of text, such as poems. The complex category is a combination of the previous two categories; the instructions in this category are the most complicated as they usually contain multiple challenging sub-instructions. While we found these categories useful to elicit a diverse instruction set, we also found that practical examples were often difficult to categorize objectively. As a result, we only used these categories for data collection, but are not reporting per-category results.
@@ 2.2 Response Evaluation Method
We adopt GPT-4o [27] to score MLLMs’ responses on each instruction and return a total score using the following prompt:
Each response is graded by first assessing how well it follows each sub-instruction, then computing the total score. Figure 6 shows an example of how responses from different MLLMs are evaluated and scored. Each sub-instruction in an instruction is assigned a maximum score ranging from 1 to 10; sum of the weight of all sub-instructions in an instruction is 10. For the example in Figure 6, there are 4 sub-instructions (denoted from S1 to S4); the first is worth 4 points and the rest is worth 2 points each. The response from GPT-4o partially follows the first sub-instruction which requires the response to be from the perspective of the dog, as the dog should not be able to see the car behind the man without turning around. The dog should be able to see the guitar, thus GPT-4o gets 2 points out of 4 for the first sub-instruction. It successfully follows the other 3 sub-instructions, achieving full score for them. Thus, the final score GPT-4o reaches is 8 out of 10. We always assign larger weight (6 if there are two sub-instructions, 4 if there are three or more sub-instructions) to the sub-instruction in the description category unless this category is absent in some cases, as usually a major part of the response is addressing this sub-instruction. For each MLLM, we compute the average score it gets on all 400 responses, and represent the ratio of the average score divided by 10 using percentage. We also compute the average score for each instruction category.
Figure 6: An example with responses from four MLLMs and their evaluation scores.
@@ 3.1 Benchmark Results
Open Source
Fuyu-8b [36] 24.52 52.06 24.52 17.06 17.18 36.43 22.62 66.67 33.09
Kosmos-2 [37] 26.06 50.95 38.52 11.55 19.78 28.70 17.26 50.83 41.88
InstructBLIP-13b [5] 38.16 50.54 39.57 29.34 38.43 42.28 12.50 50.00 30.42
Sphinx [9] 50.99 75.33 53.51 60.45 48.28 57.75 47.41 70.00 61.04
Idefics-2-8b [38] 51.42 59.37 62.73 48.07 64.09 46.20 46.51 48.33 61.97
Yi-VL-34b [39] 53.90 74.89 52.05 59.09 55.91 57.25 54.17 41.85 70.09
mPLUG-Owl2 [40] 57.86 75.01 65.25 63.39 60.26 57.70 57.22 65.00 62.08
CogVLM-Chat [8] 58.95 60.42 57.86 67.94 60.55 62.92 36.67 60.83 61.87
ShareGPT4V [41] 59.41 81.08 63.49 63.88 58.46 62.49 52.98 82.50 72.29
DeepSeek-VL-7b-chat [42] 60.96 86.31 63.26 72.11 54.79 63.75 67.39 74.17 77.85
LLaVA-1.5-7b [4] 62.18 78.00 68.60 63.95 64.18 65.89 47.31 86.67 60.75
LLaVA-NeXT-7b-vicuna [30] 62.27 79.21 68.01 65.63 60.95 63.33 46.67 90.00 65.54
Qwen-VL-Chat [7] 63.09 80.51 74.22 66.95 63.11 63.01 45.00 75.83 66.01
LLaVA-1.5-13b [4] 63.55 80.98 70.15 64.54 59.30 67.42 45.11 69.17 76.28
XComposer2-7b [43] 67.71 83.47 76.16 73.66 67.69 67.01 48.61 77.50 68.06
LLaVA-NeXT-13b-vicuna [30] 69.16 86.75 69.88 82.07 64.77 74.99 48.56 77.50 75.83
CogVLM2 [8] 73.43 87.60 74.52 83.47 71.97 77.01 71.53 90.83 87.16
InternVL-Chat-v1.5 [2] 75.42 89.13 78.21 79.92 78.16 77.54 76.11 87.50 80.92
LLaVA-NeXT-34b [30] 75.61 88.02 83.50 86.58 71.57 75.83 68.06 87.50 80.26
Phi-3-vision [31] 76.02 84.90 84.46 86.52 67.93 74.70 78.16 74.17 83.96
MiniCPM-Llama3-v2.5 [44] 76.27 84.12 79.44 80.33 81.25 76.99 64.08 81.67 76.59
LLaVA-NeXT-110b [30] 79.84 86.99 84.86 82.49 79.04 80.10 71.94 80.83 75.45
@@
Gemini-1.0-Pro [10] 70.63 82.77 72.83 78.76 76.91 71.67 81.45 89.29 84.11
Reka-Core [29] 76.95 91.05 79.91 85.16 78.98 82.08 82.53 77.50 81.08
Claude-3-Haiku [28] 78.25 86.86 77.53 90.27 73.41 82.62 82.22 57.50 86.49
Claude-3-Sonnet [28] 79.44 88.06 82.71 90.54 79.60 82.05 82.22 76.67 84.43
Claude-3-Opus [28] 84.50 90.50 86.03 91.19 83.82 85.49 85.92 65.00 86.84
GPT-4v [1] 86.11 90.03 87.61 94.59 80.12 89.37 85.63 59.17 85.26
GPT-4o [27] 88.58 90.82 92.73 94.29 85.70 90.66 87.07 92.50 86.54
@@
  • Overall, the best performance was achieved by GPT-4o [27], with a score 88.58, showcasing its superiority across different categories of instruction adherence.
  • The ability to describe content accurately was best exhibited by Reka [29]. Other models like Claude-3-Opus [28], GPT-4v [1] and GPT-4o also achieved scores higher than 90. This suggests that these models are good at generating coherent and contextually appropriate text.
  • In the genre category, the highest proficiency was shown by GPT-4v and GPT-4o with scores above 94, suggesting an exceptional grasp of language nuances. Among open-source models, Phi-3-Vision [31] and LLaVA-NeXT-34b [30] show strong performance with scores of 86.52 and 86.58, respectively. The lowest score on this metric was by Kosmos-2 [37], with a mere 11.55, pointing to difficulties in understanding or generating linguistically complex sentences.
  • GPT-4o excelled in grammar with a score of 85.70, which indicates superior ability in syntax correctness and sentence structuring that matches specific requirements in the instruction. Among the open-source models, MiniCPM-Llama3-V-2.5 [44] is notable with a score of 81.25. Contrastingly, Fuyu-8b [36] scored the lowest with 17.18, reflecting major challenges in grammar adherence.
  • GPT-4o also showed the best performance with a score of 92.73 in respecting prescribed length limits, which is crucial for tasks requiring concise and precise answers. Among open-source models, LLaVA-NeXT-110b [30] stands out with a score of 84.86.
@@
GPT-4v [1] 1926.6 56.8 77/74.4 67.6 46.5 49.9 2 86.11 1
Gemini-1.0-Pro [10] 1933.4 47.9 73.6/74.3 64.3 45.2 45.2 4 70.63 5
Claude-3-Opus [28] 1586.8 59.4 63.3/59.2 58.1 37.8 50.5 5 84.50 2
InternVL-Chat-V1-5 [2] 2187.8 45.2 82.2/82 62.8 49.3 53.5 1 75.42 4
LLaVA-NeXT-34b [23] 2028 51.1 81.1/79 48.9 47.6 47.7 3 75.61 3
@@
Correlation with other benchmarks. In Table 2, we compare the ranking of 5 state-of-the-art MLLMs on MIA-Bench as well as their meta ranking on MME [15], MMMU [3] , MMBench [45], MMVet [19], HallusionBench [46], and MathVista [18] (meta ranking is computed by averaging rankings across these benchmarks). Our findings reveal a discrepancy between the two sets of rankings. Notably, InternVL-Chat-V1.5 [2], which holds the highest meta-ranking among the five MLLMs on the other benchmarks, ranks the lowest on MIA-Bench. Conversely, Claude-3-Opus, which has the lowest meta-ranking, secures the second position on the MIA-Bench. This indicates that excelling in tasks evaluated by existing benchmarks does not necessarily translate to superior instruction adherence capability assessed by MIA-Bench.
Correlation with LLM backbone performance. To determine if the performance on MIA-Bench is attributable solely to the underlying LLMs, we also evaluate several MLLMs on IFEval [26], a benchmark that assesses the instruction adherence capability of LLMs, and compare their ranking with that on MIA-Bench. This comparison is shown in Appendix, which shows that the instruction adherence capabilities of MLLMs do not consistently align with their LLMs’ adherence capability.

### v5

@@ Abstract
We introduce MIA-Bench, a new benchmark designed to evaluate multimodal large language models (MLLMs) on their ability to strictly adhere to complex instructions. Our benchmark comprises a diverse set of 400 image-prompt pairs, each crafted to challenge the models’ compliance with layered instructions in generating accurate responses that satisfy specific requested patterns. Evaluation results from a wide array of state-of-the-art MLLMs reveal significant variations in performance, highlighting areas for improvement in instruction fidelity. Additionally, we create extra training data and explore supervised fine-tuning to enhance the models’ ability to strictly follow instructions without compromising performance on other tasks. We hope this benchmark not only serves as a tool for measuring MLLM adherence to instructions, but also guides future developments in MLLM training methods. [^1] Benchmark data and evaluation code: https://github.com/apple/ml-mia-bench.
Figure 1: An example from MIA-Bench, featuring an image and a complex instruction to test models’ compliance with layered instructions that are compositional in nature. Responses from GPT-4v (Achiam et al., 2023) and InternVL-v1.5 (Chen et al., 2024b) are evaluated using GPT-4o as the judge.
@@ 1 Introduction
Figure 2: Comparison of various multimodal LLM benchmarks. (Left) Fixed-form visual question answering, often features short answers or multi-choice formats, such as MMMU (Yue et al., 2023). This format is popular due to its ease of evaluation. (Middle) Open-ended responses, such as LLaVA-Bench (in the Wild) (Liu et al., 2023c). (Right) The proposed MIA-Bench, which also uses open-ended responses but focuses on evaluating precise adherence to complex instructions within the prompt.
The rapid advancement of Multimodal Large Language Models (MLLMs) (Achiam et al., 2023; Liu et al., 2023c; Dai et al., 2023; Liu et al., 2023b; Bai et al., 2023; Wang et al., 2023a; Lin et al., 2023b; Team, 2023; McKinzie et al., 2024) has been a defining feature of recent AI research, showcasing increased model capabilities to comprehend and respond to visual inputs, often termed as multimodal “instruction following”.
To measure the progress of instruction following, many multimodal benchmarks have been developed, which can be roughly divided into two broad categories: ($i$) fixed-form visual question answering (VQA), often with short answers or using a multi-choice QA format; and ($ii$) free-form conversations with open-ended responses. Many current benchmarks have adopted the first format, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019b), ScienceQA (Lu et al., 2022), MME (Fu et al., 2023a), MMBench (Liu et al., 2023d), SEED-Bench (Li et al., 2023a), MathVista (Lu et al., 2023), and MMMU (Yue et al., 2023). These benchmarks are popular due to their ease of use in evaluating metrics and presenting model comparisons.
However, as visual assistant models, the ability to engage users in free-form conversations is also crucial. Benchmarks in this format include LLaVA-Bench (Liu et al., 2023c), MM-Vet (Yu et al., 2023b), VisIT-Bench (Bitton et al., 2023), InfiMM-Eval (Han et al., 2023), and the most recent Vibe-Eval (Padlewski et al., 2024) and LLaVA-Bench-Wilder (Li et al., 2024a). Typically, the free-form model responses are evaluated using external models as the judge. These benchmarks are closer to daily-life visual chat scenarios; however, the type of “instruction following” examined in these benchmarks usually gauges a model’s ability to perform tasks in a broad, often loosely defined manner. Yet, the precise adherence to complex instructions within prompts – a critical aspect for evaluating LLMs (Chia et al., 2023; Zhou et al., 2023; Qin et al., 2024) – remains less explored in the context of multimodal LLMs.
@@
We evaluate a wide array of MLLMs on the proposed benchmark, ranging from closed-source models (e.g., GPT-4o (OpenAI, 2024), Gemini Pro (Team, 2023), Claude-3 (Anthropic, 2024), Reka (Team et al., 2024)) to open-source ones (e.g., LLaVA-NeXT (Liu et al., 2024a), Intern-VL-Chat-1.5 (Chen et al., 2024b), CogVLM2 (Wang et al., 2023a), Phi-3-Vision (Abdin et al., 2024)). Our investigations reveal notable variations in model performance, highlighting great opportunities for improvement.
@@
Our contributions are summarized as follows. ($i$) We construct MIA-Bench, a new benchmark to comprehensively evaluate MLLMs on their capability to strictly adhere to instructions. ($ii$) We provide a detailed analysis of popular MLLMs, and suggest training methods for enhanced instruction following. For this purpose, we created training data and conducted experiments for additional supervised fine-tuning. We hope this benchmark can serve as a useful resource to stimulate further research on multimodal instruction adherence.
@@ 2 MIA-Bench
MIA-Bench consists of 400 image-prompt pairs, with examples shown in Figure 3. The images are collected from diverse sources, including COCO 2017 validation set (Lin et al., 2015), SBU (Ordonez et al., 2011), TextVQA (Singh et al., 2019a), and Flickr. Images in the Flickr subset are photos of a variety of themes, including animals, art, architectures, text, food, math, etc. Images from the other three sources are randomly sampled from each corresponding source. Figure 4 shows the top 15 image content categories and the distribution of the 8 sub-instruction categories in MIA-Bench. The image content is labeled using GPT-4v. For each image, we manually write diverse and challenging instructions that contain multiple sub-instructions.
@@
  • Image-dependent. MMStar (Chen et al., 2024a) pointed out that on some multimodal benchmarks, MLLMs can generate correct answers without accessing images half of the time. Multi-modal capabilities are necessary to correctly answer MIA-Bench prompts.
@@ 2.1 Instruction Categories
Figure 5 shows the most frequently used verbs and co-occurring nouns in MIA-Bench. To guarantee the diversity of prompts, when writing the instructions, we construct instructions of various levels of complexity: basic, intermediate, advanced, creative, and complex.
@@
The basic category is the simplest; the instructions normally only contain one or two sub-instructions, such as “What is the color of the cat?”, or “Describe the sofa in two words.”. The intermediate category consists of instructions that contain three or more sub-instructions, but are in general easy for MLLMs to follow. The advanced category contains instructions that are challenging and contain three or more sub-instructions. The creative category contains instructions that instruct MLLMs to generate creative pieces of text, such as poems. The complex category is a combination of the previous two categories; the instructions in this category are the most complicated as they usually contain multiple challenging sub-instructions. While we found these categories useful to elicit a diverse instruction set, we also found that practical examples were often difficult to categorize objectively. As a result, we only used these categories for data collection, but are not reporting per-category results.
@@ 2.2 Response Evaluation Method
To automatically evaluate MLLMs’ performance with the proposed MIA-Bench at scale, we adopt GPT-4o (OpenAI, 2024) as a judge model to score MLLMs’ responses on each instruction and return a total score based on different criteria mentioned above. Specifically, we design the following text prompt template:
Each response is graded through a systematic process that begins by thoroughly evaluating how well it adheres to each specific sub-instruction provided within the overall task. This initial assessment focuses on the degree of compliance and alignment with the requirements outlined in the sub-instructions. After this detailed evaluation of individual components, a total score is computed by aggregating the performance across all sub-instructions. Each sub-instruction within the broader instruction is assigned a maximum possible score, which can range from 1 to 10. The specific score assigned to a sub-instruction reflects its relative significance and difficulty within the context of the overall task. To ensure fairness and consistency, the total weight of all sub-instructions for any given instruction is set to sum up to 10. This carefully calibrated scoring system has been meticulously designed to capture and reflect the varying levels of complexity and importance associated with each aspect of the task. By doing so, it ensures that more critical or challenging elements are appropriately weighted and that the evaluation process provides a comprehensive and accurate representation of the response’s quality.
Figure 6 shows an example of how responses from different MLLMs are evaluated and scored. For example, length limits are often binary in nature—either met or not met—hence the single point allocation. In contrast, a description task may require the model to handle multiple layers of complexity, including accuracy, detail, and relevance, which justified a higher score. For the example in Figure 6, there are 4 sub-instructions (denoted from S1 to S4); the first is worth 4 points and the rest is worth 2 points each. The response from GPT-4o partially follows the first sub-instruction which requires the response to be from the perspective of the dog, as the dog should not be able to see the car behind the man without turning around. The dog should be able to see the guitar, thus GPT-4o gets 2 points out of 4 for the first sub-instruction. It successfully follows the other 3 sub-instructions, achieving full score for them. Thus, the final score GPT-4o reaches is 8 out of 10. We always assign larger weight (6 if there are two sub-instructions, 4 if there are three or more sub-instructions) to the sub-instruction in the description category unless this category is absent in some cases, as usually a major part of the response is addressing this sub-instruction. For each MLLM, we compute the average score it gets on all 400 responses, and represent the ratio of the average score divided by 10 using percentage. We also compute the average score for each instruction category.
Figure 6: An example featuring responses from four MLLMs along with their evaluation scores, presented in a point-by-point format.
@@ 3.1 Benchmark Results
Open Source 1b-8b
Fuyu-8b (Bavishi et al., 2023) 24.52 52.06 24.52 17.06 17.18 36.43 22.62 66.67 33.09
Kosmos-2 (Peng et al., 2023a) 26.06 50.95 38.52 11.55 19.78 28.70 17.26 50.83 41.88
Sphinx (Lin et al., 2023b) 50.99 75.33 53.51 60.45 48.28 57.75 47.41 70.00 61.04
Idefics-2-8b (Laurençon et al., 2024) 51.42 59.37 62.73 48.07 64.09 46.20 46.51 48.33 61.97
mPLUG-Owl2 (Ye et al., 2023b) 57.86 75.01 65.25 63.39 60.26 57.70 57.22 65.00 62.08
CogVLM-Chat (Wang et al., 2023a) 58.95 60.42 57.86 67.94 60.55 62.92 36.67 60.83 61.87
ShareGPT4V (Chen et al., 2023b) 59.41 81.08 63.49 63.88 58.46 62.49 52.98 82.50 72.29
DeepSeek-VL-7b-chat (Lu et al., 2024) 60.96 86.31 63.26 72.11 54.79 63.75 67.39 74.17 77.85
LLaVA-1.5-7b (Liu et al., 2023c) 62.18 78.00 68.60 63.95 64.18 65.89 47.31 86.67 60.75
LLaVA-NeXT-7b-vicuna (Liu et al., 2024a) 62.27 79.21 68.01 65.63 60.95 63.33 46.67 90.00 65.54
Qwen-VL-Chat (Bai et al., 2023) 63.09 80.51 74.22 66.95 63.11 63.01 45.00 75.83 66.01
XComposer2-7b (Dong et al., 2024a) 67.71 83.47 76.16 73.66 67.69 67.01 48.61 77.50 68.06
CogVLM2 (Wang et al., 2023a) 73.43 87.60 74.52 83.47 71.97 77.01 71.53 90.83 87.16
Phi-3-vision (Abdin et al., 2024) 76.02 84.90 84.46 86.52 67.93 74.70 78.16 74.17 83.96
MiniCPM-Llama3-v2.5 (Hu et al., 2023) 76.27 84.12 79.44 80.33 81.25 76.99 64.08 81.67 76.59
Open Source 8b-13b
InstructBLIP-13b (Dai et al., 2023) 38.16 50.54 39.57 29.34 38.43 42.28 12.50 50.00 30.42
LLaVA-1.5-13b (Liu et al., 2023c) 63.55 80.98 70.15 64.54 59.30 67.42 45.11 69.17 76.28
LLaVA-NeXT-13b-vicuna (Liu et al., 2024a) 69.16 86.75 69.88 82.07 64.77 74.99 48.56 77.50 75.83
Open Source 13b-110b
Yi-VL-34b (AI et al., 2024) 53.90 74.89 52.05 59.09 55.91 57.25 54.17 41.85 70.09
InternVL-Chat-v1.5 (Chen et al., 2024b) 75.42 89.13 78.21 79.92 78.16 77.54 76.11 87.50 80.92
LLaVA-NeXT-34b (Liu et al., 2024a) 75.61 88.02 83.50 86.58 71.57 75.83 68.06 87.50 80.26
LLaVA-NeXT-110b (Liu et al., 2024a) 79.84 86.99 84.86 82.49 79.04 80.10 71.94 80.83 75.45
@@
Gemini-1.0-Pro (Team, 2023) 70.63 82.77 72.83 78.76 76.91 71.67 81.45 89.29 84.11
Reka-Core (Team et al., 2024) 76.95 91.05 79.91 85.16 78.98 82.08 82.53 77.50 81.08
Claude-3-Haiku (Anthropic, 2024) 78.25 86.86 77.53 90.27 73.41 82.62 82.22 57.50 86.49
Claude-3-Sonnet (Anthropic, 2024) 79.44 88.06 82.71 90.54 79.60 82.05 82.22 76.67 84.43
Claude-3-Opus (Anthropic, 2024) 84.50 90.50 86.03 91.19 83.82 85.49 85.92 65.00 86.84
GPT-4v (Achiam et al., 2023) 86.11 90.03 87.61 94.59 80.12 89.37 85.63 59.17 85.26
GPT-4o (OpenAI, 2024) 88.58 90.82 92.73 94.29 85.70 90.66 87.07 92.50 86.54
@@
  • Overall, the best performance was achieved by GPT-4o (OpenAI, 2024), with a score 88.58, showcasing its superiority across different categories of instruction adherence.
  • The ability to describe content accurately was best exhibited by Reka (Team et al., 2024). Other models like Claude-3-Opus (Anthropic, 2024), GPT-4v (Achiam et al., 2023) and GPT-4o also achieved scores higher than 90. This suggests that these models are good at generating coherent and contextually appropriate text.
  • In the genre category, the highest proficiency was shown by GPT-4v and GPT-4o with scores above 94, suggesting an exceptional grasp of language nuances. Among open-source models, Phi-3-Vision (Abdin et al., 2024) and LLaVA-NeXT-34b (Liu et al., 2024a) show strong performance with scores of 86.52 and 86.58, respectively. The lowest score on this metric was by Kosmos-2 (Peng et al., 2023a), with a mere 11.55, pointing to difficulties in understanding or generating linguistically complex sentences.
  • GPT-4o excelled in grammar with a score of 85.70, which indicates superior ability in syntax correctness and sentence structuring that matches specific requirements in the instruction. Among the open-source models, MiniCPM-Llama3-V-2.5 (Hu et al., 2023) is notable with a score of 81.25. Contrastingly, Fuyu-8b (Bavishi et al., 2023) scored the lowest with 17.18, reflecting major challenges in grammar adherence.
  • GPT-4o also showed the best performance with a score of 92.73 in respecting prescribed length limits, which is crucial for tasks requiring concise and precise answers. Among open-source models, LLaVA-NeXT-110b (Liu et al., 2024a) stands out with a score of 84.86.
@@
GPT-4o (OpenAI, 2024) 2328.7 69.1 83.3/82.1 66.5 67.5 63.8 1 88.58 1
GPT-4v (Achiam et al., 2023) 1926.6 56.8 77/74.4 67.6 46.5 49.9 3 86.11 2
Gemini-1.0-Pro (Team, 2023) 1933.4 47.9 73.6/74.3 64.3 45.2 45.2 5* 70.63 6
Claude-3-Opus (Anthropic, 2024) 1586.8 59.4 63.3/59.2 58.1 37.8 50.5 5* 84.50 3
InternVL-Chat-V1-5 (Chen et al., 2024b) 2187.8 45.2 82.2/82 62.8 49.3 53.5 2 75.42 5
LLaVA-NeXT-34b (Li et al., 2024a) 2028 51.1 81.1/79 48.9 47.6 47.7 4 75.61 4
@@
Correlation with other benchmarks. In Table 2, we compare the ranking of 5 state-of-the-art MLLMs on MIA-Bench as well as their meta ranking on MME (Fu et al., 2023a), MMMU (Yue et al., 2023) , MMBench (Liu et al., 2024b), MMVet (Yu et al., 2023b), HallusionBench (Guan et al., 2023), and MathVista (Lu et al., 2023) (meta ranking is computed by averaging rankings across these benchmarks). Our findings reveal a discrepancy between the two sets of rankings. Notably, InternVL-Chat-V1.5 (Chen et al., 2024b), which holds the highest meta-ranking among the five MLLMs on the other benchmarks, ranks the lowest on MIA-Bench. Conversely, Claude-3-Opus, which has the lowest meta-ranking, secures the second position on the MIA-Bench. This indicates that excelling in tasks evaluated by existing benchmarks does not necessarily translate to superior instruction adherence capability assessed by MIA-Bench.
Figure 7: Study of correlation between scores on MIA-Bench and IFEval.
Correlation with LLM backbone performance. To determine if the performance on MIA-Bench is attributable solely to the underlying LLMs, we also evaluate several MLLMs on IFEval (Qin et al., 2024), a benchmark that assesses the instruction adherence capability of LLMs, and compare their ranking with that on MIA-Bench. This comparison is shown in Figure 7, which shows that the instruction adherence capabilities of MLLMs do not consistently align with their LLMs’ adherence capability.
@@
We further extend our study by evaluating the models using various versions of ChatGPT as judge models, including gpt-4o-mini-2024-07-18, gpt-4o-2024-05-13, gpt-4o-2024-11-20, and chatgpt-4o-latest, and show the scores and ranking in Table 3 with detailed results in Table 678 and 9. We find that the ranking is consistent even though the judge models are different.
Model chatgpt-4o-latest gpt-4o-2024-11-20 gpt-4o-2024-05-13 gpt-4o-mini-2024-07-18
Score Ranking Score Ranking Score Ranking Score Ranking
GPT-4o 89.69 1 89.94 1 90.97 1 81.36 1
Claude-3-Opus 86.16 2 84.89 2 85.61 2 78.95 2
Reka 83.09 3 82.68 3 83.99 3 77.70 3
MiniCPM-Llama3-V2.5 78.10 4 78.75 4 79.80 4 73.72 4
Gemini 75.77 5 76.32 5 77.36 5 67.45 5
LLaVA-1.5-13b 66.78 6 66.05 6 68.39 7 61.54 6
ShareGPT4v 66.61 7 65.72 7 68.90 6 60.30 7
Idefics-2-8b 53.51 8 53.61 8 54.18 8 44.28 8
Table 3: Comparison of scores and rankings across different judge models. The ranking is stable.